Describe and compute the takt time based on the information below.
Operating Time:
The plant worked one shift of eight hours, with two 15 minute breaks and one 15 minute stand up department meeting.
Customer Demand:
Part A Quantity 40, Part B Quantity 20, Part C Quantity 50, and Part D Quantity 30
Answer #1

Takt time means that maximum time needed to produce one unit of product inorder to meet the customer demand .

Takt time =

Given that one shift of eight hours = 60min /hr * 8 hour

= 480 mints /shift

Break = 2*15mint /shift

= 30 mint /shift

Meeting = 15 mint /shift

Net operating time = 480 -30 -15

= 480 - 45

= 435 mints /shift

Takt time = 435 / quantity

Part A

Takt time = 435/40

=10.875 minute /unit

Part B

Takt time = 435 /20

= 21.75 minute/ unit

Part C

Takt time = 435/ 50

= 8.7 minute / unit

Part D

Takt time = 435/30

= 14.5 minute / unit